OpenStudy (anonymous):

PLEASE HELP stuck on this stupid question for a hour i suck at math :C Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? Explain whether the graph has a maximum or minimum point. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

x^2-4x+4

OpenStudy (anonymous):

y = x^2 - 2x * Explain whether the graph has a maximum or minimum point. It has a minimum point since it opens up. * Find the vertex and x-intercepts of the graph. y = x^2 - 2x = (x - 1)^2 - 1 Vertex: (1, -1) y = x^2 - 2x = x(x - 2) x intercepts: (0, 0), (0, 2)

OpenStudy (anonymous):

nice Kdwndr

OpenStudy (anonymous):

Thank you really i suck at this topic big help!

OpenStudy (anonymous):

you don't suck ,, it's just you didn't try enough.

OpenStudy (anonymous):

no i really have no clue how to o this i read the lesson and i do not understand at all i really am bad at math it just doesn't click

OpenStudy (anonymous):

Just practice and you'll get better

OpenStudy (anonymous):

I am not good at math either, I just keep trying.

OpenStudy (anonymous):

I do try i take online classes i have no one to explain things to me so it makes it 10 times harder.

OpenStudy (anonymous):

I can understand, but don't give up! :)

OpenStudy (anonymous):

Thanks :p